OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1

OpenStudy (anonymous):

please help me ahh!

OpenStudy (anonymous):

x = -7 y = 9

OpenStudy (anonymous):

20x+16y=4.............(1) 20x+15y=-5.................(2) subtracting 2 from 1st equation,we get y=9